Shift summary for OWL shift 11/Jan/2000


Subject: Shift summary for OWL shift 11/Jan/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:17:30 CST


                     Shift summary for owl shift 11/Jan/2000

On Shift: RK, AG

Hours of beam available at all intensities: 1.8
Hours of beam available at nominal intensity: 1.8
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 1.5

Accounting of lost time:

o Accelerator problems.

o Front end problems.

========================================================================

Tapes brought to FCC this shift: none.

========================================================================

Accesses during shift: none.

========================================================================

Runs ended this shift:

15477 - KAMI_KE3
 
=======================================================================

Overview of shift:

Very poor accelerator performance. Got a little bit of KAMI_KE3 data at
the beginning of the shift and then some more at the end. Tevatron
is still in bad shape, also front end is unstable - lots of ktevana
errors.

========================================================================

Description of Shift:

run spill time comment
------------------------------------------------------------------------

15475 104 0002 Noticed that NM3S tripped off
                        -> called MCR and they fixed it.

                 0025 Stop KAMI_KE3 run 15475 to flip NM4 polarity while
                        beam is down for a while.

                 0105 Re-make L3-drivers with +BASWAP patchy flag to
                        incorporate ADC as shown at the top of p.42 (KAMI
                        Log I).

15477 0110 Start run 15477.

                 0500 Beam is back ... Not really, Quench again! Est. 1h

                 0605 Stop run 15477.

                 0615 Run 15478 is started.

           14 0632 Beam is back...

========================================================================



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:17:38 CST